# Multi-state Models For Modeling Disease Histories Based On Longitudinal Data

### v1

@@ 5.1 UK Biobank data on history of CKD stages
The UKB dataset curated by Gorski et al. [50] contains $2,080,372$ eGFR measurements from $n=454,071$ subjects, collected from two study center assessments and electronic health records (eHRs).
We apply KDIGO definitions [6] to define CKD stages: Healthy (eGFR $\geq$ 60 mL/min/1.73m2), Mild CKD (eGFR $\in[30,60)$ mL/min/1.73m2), Severe CKD (eGFR $\in[15,30)$ mL/min/1.73m2), and ESKD (eGFR << 15 mL/min/1.73m2 or dialysis or kidney transplant). Healthy, Mild and Severe CKD are transient events, while both ESKD and Death are absorbing (see Figure 1).
The final multi-state dataset contains $142,667$ subjects and $19,293$ non-censoring transitions (Table 1; see Appendix A.2.1 for further details on data preprocessing).
@@ 5.2 Model specification
We draw on the conclusions from Section 4 to specify an SSTS and an MTS PAM with interval mid-point as estimation point. Here, we choose chronological age to be the primary time scale due to the non-informativeness of the study entry time point [24]. The baseline hazard model equation of the SSTS PAM for transitions $k$ (cf. Figure 1; Equation (3)) is
@@
For IEB-related analyses (Section 5.3.2), we include the following risk factors in the model (stratified by transition; cf. Equation (6)): (i) the genetic variant rs77924615 (effect allele G); (ii) sex; (iii) a polygenic risk score (PGS) to control for other genetic variants associated with kidney function [52, 24]; (iv) binary diabetes and smoking status; and (v) body mass index (BMI) and urine albumin-to-creatinine ratio (uACR), both from UKB study center visits only.
@@ 5.3.1 Estimation of baseline hazards and state-entry times
For transitions out of Mild CKD, evaluation involves two time dimensions: age (chronological age) and $age_{1}$ (age at onset of Mild CKD), which jointly determine $t_{1}$ (time since onset of Mild CKD) via $t_{1}=age-age_{1}$. The ”Mild CKD $\rightarrow$ Severe CKD” transition probabilities are highest for subjects with early CKD onset age and, given $age_{1}$, increase over age. Direct ”Mild CKD $\rightarrow$ Death” transition probabilities again increase sharply with age (starting at around 50), independently of $age_{1}$ (Figures A.8 and A.9).
For transitions out of Severe CKD, we have an additional third time dimension $age_{2}$ (age at progression to Severe CKD), which determines $t_{2}$ (time since onset of Mild CKD) via $t_{2}=age-age_{2}$. Transition probabilities for ”Severe CKD $\rightarrow$ ESKD” rise with decreasing $age_{1}$ and $age_{2}$; that is, they are highest for early transitions into Mild and Severe CKD. Given $age_{1}$ and $age_{2}$, the transition probabilities increase with age (Figures A.10 and A.11). ”Severe CKD $\rightarrow$ Death” transitions show a similar pattern as direct ”Mild CKD $\rightarrow$ Death” transitions (Figures A.12 and A.13).
@@
Finally, we observe that the effects of state-entry times (here: age at CKD onset and progression) are mostly linear, though not always (Figures Figure 7d+e and A.14). This underlines the importance of the modeling flexibility provided by this PAM-based framework.
@@ 5.3.2 Index event bias and genetic effect estimation
G + Sex + PGS + Diabetes + Smoking + BMI + uACR 2→3 -0.159 0.178 0.371
@@
The most straightforward approach is to directly adjust the model for other risk factors. Here we include sex, PGS, diabetes, smoking, BMI, and uACR (see Section 5.2). Inclusion of these risk factors leaves the genetic effect on CKD onset largely unchanged, while substantially attenuating the genetic effect on progression to Severe CKD from 0.12-0.12 to +0.05+0.05 (Table 2). So while there is sound evidence for a hazard-increasing effect of the UMOD genetic variant on CKD onset, there seems to be neither a detrimental nor a protective effect on CKD progression. This answers the second question posed at the very beginning of this paper.
An alternative to the direct adjustment for other risk factors is to use propensity scores to compute subject- and state-specific stabilizing weights for the estimation of a weighted PAM (Appendix A.2.2). This weighting approach produces similar results as the adjustment approach regarding the estimation of rs77924615 effects on CKD onset and progression (Table A.12). However, we note that the incorporation of time-varying covariates requires the estimation of time-varying weights and, while feasible, is much more complex [53, 54, 55].
@@ 5.3.3 Evaluating the impact of interval-censoring on the CKD multi-state model estimation
Due to the dense eHRs in our dataset (median time between eGFR measurements: 1.44 years), the overall degree of interval-censoring is limited. Furthermore, death time points are not affected by interval-censoring as the corresponding information stems from exactly recorded death registries. As expected, we observe informative observation [36, 37] as interval length is related to CKD stage (median time between observations: 1.46, 0.68, and 0.27 years for Healthy, Mild CKD, and Severe CKD, respectively). This means that less healthy individuals – i.e., those who are more likely to have CKD-related events – tend to have more detailed eHR trajectories and, hence, smaller interval lengths (cf. Section 2.4).
@@ 6 Discussion
Regarding the choice of time scales, a stratified single time scale proved more robust to model misspecification in terms of baseline hazard estimation than using multiple time scales. The SSTS PAM also showed a better fit on the empirical data from UKB. Interestingly, this is also in line with the leukemia example by Iacobelli and Carstensen [16], whose final model eventually does not contain multiple time scales either. In terms of estimating the hazards of fixed effects (e.g., risk factors), both approaches were found to be robust.
@@
On UKB data, SSTS and MTS PAMs mostly produced similar estimates of CKD onset and progression hazards and transition probabilities. In particular, we found progression risk to Severe CKD to be highest for subjects with early onset of Mild CKD and to further rise with age. From there, the risk of further progressing to ESKD was highest for subjects with early onset of Mild CKD and progression to Severe CKD, again increasing with age. We also found the G allele of the genetic variant rs77924615 (UMOD locus) to increase the risk of CKD onset, in line with literature [56, 24]. Importantly, we observed an initial protective effect of this G allele on progression into Severe CKD and ESKD; yet this effect disappeared after adjustment for other risk factors of CKD and CKD progression. The initial protective effect thus seems to be attributable to IEB.
@@
First, the individual simulation setups could be expanded even further. For instance, it would be interesting to expand the IEB simulations to scenarios with dependence structures between risk factors in the healthy population. In particular, for specific application settings it could be informative to directly use empirical joint distributions of risk factors for simulations.
Second, the recovery of baseline hazards by PAMs under interval-censoring is still suboptimal. This is because direct incorporation of interval-censoring into the PEM likelihood, while maintaining its proportionality to a Poisson likelihood, remains a challenge.
Third, our proposed handling of IEB relies upon other risk factors being observable, which is often not (or only partially) the case: In our UKB dataset, for instance, the risk factors BMI and uACR are available only from study center assessments, but not from eHRs. And while off-the-shelf bias correction methods for linear and binary outcomes exist [57, 58, 59], their applicability to multi-state models is still to be explored.
Finally, as UKB participants are rather healthy [60], CKD progressions in this UKB dataset are relatively rare. This, in turn, makes the estimation of (especially small-scale, genetic) risk factors more difficult. It would thus be interesting to conduct similar CKD onset and progression analyses in less healthy cohorts.
@@
Multi-state PAMs (including all necessary preprocessing) are implemented in the R package pammtools [19], which uses mgcv::gam()/bam() [45] for efficient and flexible estimation. All code required for running the simulations (Section 4), for the UK Biobank analyses (Section 5), as well as for creating the corresponding figures and tables, is available from the GitHub repository https://github.com/survival-org/msm4diseaseHistories.
@@ A.1.1 Transformation to piecewise exponential data
Note that accounting for left-truncation, time-varying covariates, and time-varying effects all require similar data transformations and can therefore seamlessly be incorporated into this data preprocessing step [11]. To capture changes in time-varying covariates accurately, additional time cut points can be added at times where the time-varying covariates change [19].
PEMs then use $d_{ij}\sim Po(\mu_{ij})$ as outcome of a Poisson regression model with expected conditional event rate $\mu_{ij}=h_{ij}y_{ij}$ and logarithmized time-at-risk entering as offset $o_{ij}$, to estimate the hazard rate $h_{ij}$ as the product of a piecewise constant baseline hazard rate (for each interval) and the exponential of a linear predictor. PEMs can be shown to be equivalent to the Cox Proportional Hazards model under certain assumptions [61], but since they do not rely on the non-parametric estimation of the baseline hazard (due to assuming piecewise-constant hazards), they are fully parametric with respect to parameter estimation. Still, we here refer to PEMs as semi-parametric approach, analogously to the Cox model, because no parametric event time distribution is assumed - in contrast to parametric approaches such as Accelerated Failure Time models [62].
@@ A.2 Additional information on CKD onset and progression analyses in UK Biobank
A.2.1 Further details on the UK Biobank dataset
Based on the UKB dataset curated by Gorski et al. [50], we include unrelated UKB participants of European ancestry (in line with Wiegrebe et al. [24]) with available eHR data. We exclude eGFR measurements at age <35<35 years as well as between six months before and after acute kidney injury and nephrectomy, as in Wiegrebe et al. [24]. To account for natural fluctuations when measuring eGFR, we require two subsequent eGFR measurements to fall below the respective transition thresholds for transitions into Mild and Severe CKD. For computational reasons, we round subject age to years. We impute missing interim CKD stages for ”jump” transitions ($0\rightarrow 2$, $0\rightarrow 3$, and $1\rightarrow 3$) – which in the case of CKD are due to interval-censoring and not to some separate transition mechanism – by linearly interpolating eGFR values. (This only applies to 213 subjects.)
A.2.2 Propensity score-based correction for index event bias
@@ A.4 Supplementary tables
Table A.3: Mean overage of fixed effect log-hazards estimated using SSTS and MTS PAMs on data simulated from SSTS and MTS DGPs.

### v2

@@ 5.1 UK Biobank data on history of CKD stages
The UKB dataset curated by Gorski et al. [Gorski et al., 2025] contains $2,080,372$ eGFR measurements from $n=454,071$ subjects, collected from serum creatinine measurements from blood drawn at two study center assessments as well as from electronic health records (eHRs). We include unrelated UKB participants of European ancestry (in line with Wiegrebe et al. [Wiegrebe et al., 2024]) with available eHR data. We exclude eGFR measurements at age <35<35 years as well as between six months before and after acute kidney injury and nephrectomy, as in Wiegrebe et al. [Wiegrebe et al., 2024].
We apply KDIGO definitions [Group, 2013] to define CKD stages: Healthy (eGFR $\geq$ 60 mL/min/1.73m2), Mild CKD (eGFR $\in[30,60)$ mL/min/1.73m2), Severe CKD (eGFR $\in[15,30)$ mL/min/1.73m2), and ESKD (eGFR << 15 mL/min/1.73m2 or dialysis or kidney transplant). Healthy, Mild and Severe CKD are transient events, while both ESKD and Death are absorbing (see Figure 1); the Death state is defined based on available death registries linked to the UK Biobank, which provide the exact date of death (along with primary and contributory causes) according to the ICD-10 system.
To account for natural fluctuations when measuring eGFR, we require two subsequent eGFR measurements to fall below the respective transition thresholds for transitions into Mild and Severe CKD. For computational reasons, we round subject age to years. We impute missing interim CKD stages for ”jump” transitions ($0\rightarrow 2$, $0\rightarrow 3$, and $1\rightarrow 3$) – which in the case of CKD are due to interval-censoring and not to some separate transition mechanism – by linearly interpolating eGFR values. (This only applies to 213 subjects.)
The final multi-state dataset contains $142,667$ subjects and $19,293$ non-censoring transitions (Table 1).
@@ 5.2 Model specification
We draw on the conclusions from Section 4 to specify an SSTS and an MTS PAM with interval mid-point as estimation point. Here, we choose chronological age to be the primary time scale due to the non-informativeness of the study entry time point [Wiegrebe et al., 2024]. The baseline hazard model equation of the SSTS PAM for transitions $k$ (cf. Figure 1; Equation (4)) is
@@
For IEB-related analyses (Section 5.3.2), we include the following risk factors in the model (stratified by transition; cf. Equation (7)): (i) the genetic variant rs77924615 (effect allele G); (ii) sex; (iii) a polygenic risk score (PGS) to control for other genetic variants associated with kidney function [Stanzick et al., 2023; Wiegrebe et al., 2024]; (iv) binary diabetes and smoking status; and (v) body mass index (BMI) and urine albumin-to-creatinine ratio (uACR), both from UKB study center visits only.
@@ 5.3.1 Estimation of baseline hazards and state-entry times
For transitions out of Mild CKD, evaluation does not only depend on the single time dimension of chronological age, but also on $age_{1}$ (age at onset of Mild CKD) and/or $t_{1}$ (time since onset of Mild CKD), with $age=age_{1}+t_{1}$; see subjects B and C of 2 with $age_{1}=age_{onset}$ and $t_{1}=time_{onset}$. The ”Mild CKD $\rightarrow$ Severe CKD” transition probabilities are highest for subjects with early CKD onset (i.e., small $age_{1}$) and, for a given $age_{1}$, increase over age. Direct ”Mild CKD $\rightarrow$ Death” transition probabilities again increase sharply with age (starting at around 50), independently of $age_{1}$ (Figures A.8 and A.9).
For transitions out of Severe CKD, we have additional dependence upon $age_{2}$ (age at progression to Severe CKD) and/or $t_{2}$ (time since onset of Mild CKD), with $age=age_{2}+t_{2}$; see subject B of 2 with $age_{2}=age_{progression}$ and $t_{2}=time_{progression}$. Transition probabilities for ”Severe CKD $\rightarrow$ ESKD” rise with decreasing $age_{1}$ and $age_{2}$; that is, they are highest for subjects with early transitions into Mild and Severe CKD. Given $age_{1}$ and $age_{2}$, the transition probabilities increase with age (Figures A.10 and A.11). ”Severe CKD $\rightarrow$ Death” transitions show a similar pattern as direct ”Mild CKD $\rightarrow$ Death” transitions (Figures A.12 and A.13).
@@
Finally, we observe that the effects of state-entry times (here: age at CKD onset and progression) are mostly linear, though not always (Figures 7d+e and A.14). This underlines the importance of the modeling flexibility provided by this PAM-based framework.
@@ 5.3.2 Index event bias and genetic effect estimation
G + Sex + PGS + Diabetes + Smoking + BMI + uACR (AIC: 175,181) 2→3 -0.159 0.178 0.371
@@
The most straightforward approach is to directly adjust the model for other risk factors, under the assumption that all relevant other risk factors are known and observable and, thus, no unmeasured confounding remains after adjusting for them. Here we include sex, PGS, diabetes, smoking, BMI, and uACR (see Section 5.2). Inclusion of these risk factors leaves the genetic effect on CKD onset largely unchanged, while substantially attenuating the genetic effect on progression to Severe CKD from 0.12-0.12 to +0.05+0.05 (Table 2). Additionally, the stepwise inclusion of the above risk factors also improves the model fit in terms of the Akaike Information Criterion. So while there is sound evidence for a hazard-increasing effect of the UMOD genetic variant on CKD onset, there seems to be neither a detrimental nor a protective effect on CKD progression. This answers the second question posed at the very beginning of this paper.
An alternative to the direct adjustment for other risk factors is to use propensity scores to compute subject- and state-specific stabilizing weights for the estimation of a weighted PAM (Appendix A.2.1). This weighting approach produces similar results as the adjustment approach regarding the estimation of rs77924615 effects on CKD onset and progression (Table A.12). However, we note that the incorporation of time-varying covariates requires the estimation of time-varying weights and, while feasible, is much more complex [Hernán et al., 2000, 2009; Hernán and Robins, 2020].
@@ 5.3.3 Evaluating the impact of interval-censoring on the CKD multi-state model estimation
Due to the dense eHRs in our dataset (median time between eGFR measurements: 1.44 years), the overall degree of interval-censoring is limited. Furthermore, death time points are not affected by interval-censoring as the corresponding information stems from exactly recorded death registries. As expected, we observe informative observation [Sperrin et al., 2017; Bible et al., 2024] as interval length is related to CKD stage (median time between observations: 1.46, 0.68, and 0.27 years for Healthy, Mild CKD, and Severe CKD, respectively). This means that less healthy individuals – i.e., those who are more likely to have CKD-related events – tend to have more detailed eHR trajectories and, hence, smaller interval lengths (cf. Section 2.4). As suggested in the literature [cf. Cook and Lawless, 2018, Section 5.4], Poisson processes can be used to handle informative observation – and our PAM-based framework is inherently linked to a Poisson process via likelihood equivalence. In addition, the flexibility of our non-Markov framework allows for further mitigation of any potential biases from informative observation via direct incorporation of (history-dependent, time-varying) characteristics of the observation process into the model – for example in the form of the number of visits or mean inter-visit time.
@@ 6 Discussion
Regarding the choice of time scales, a stratified single time scale proved more robust to model misspecification in terms of baseline hazard estimation than using multiple time scales. The SSTS PAM also showed a better fit on the empirical data from UKB. Interestingly, this is also in line with the leukemia example by Iacobelli and Carstensen [Iacobelli and Carstensen, 2013], whose final model eventually does not contain multiple time scales either. Therefore, a relatively simple (single time scale) model appears to suffice even for estimation of hazards from complex DGPs with multiple time scales. In terms of estimating the hazards of fixed effects (e.g., risk factors), both approaches were found to be robust.
@@
On UKB data, SSTS and MTS PAMs mostly produced similar estimates of CKD onset and progression hazards and transition probabilities. In particular, we found progression risk to Severe CKD to be highest for subjects with early onset of Mild CKD and to further rise with age. From there, the risk of further progressing to ESKD was highest for subjects with early onset of Mild CKD and progression to Severe CKD, again increasing with age. We also found the G allele of the genetic variant rs77924615 (UMOD locus) to increase the risk of CKD onset, in line with literature [Devuyst and Pattaro, 2018; Wiegrebe et al., 2024]. Importantly, we observed an initial protective effect of this G allele on progression into Severe CKD and ESKD; yet this effect disappeared after adjustment for other risk factors of CKD and CKD progression. The initial protective effect thus seems to be attributable to IEB.
@@
First, the individual simulation setups could be expanded even further. The simulations regarding choice of time scales and IEB, for instance, could be enhanced by also considering other DGPs beyond the piecewise exponential distribution – though we do note that the latter one is particularly flexible and that the simulations on interval-censoring underlined the ability of PAMs to adequately estimate multiple distinct baseline hazards. In addition, it would be interesting to expand the IEB simulations to scenarios with dependence structures between risk factors in the healthy population. In particular, for specific application settings it could be informative to directly use empirical joint distributions of risk factors for simulations.
Second, the recovery of baseline hazards by PAMs under interval-censoring is still suboptimal. This is because direct incorporation of interval-censoring into the PEM likelihood, while maintaining its proportionality to a Poisson likelihood, remains a challenge. In this context, it would also be of interest to conduct further simulations regarding the impact of interval-censoring within complex multi-state model settings, because interval-censored transition times out of the initial state also induce censoring in the observed state-entry times of subsequent states.
Third, our proposed handling of IEB relies upon other risk factors being observable and on the assumption of no unmeasured confounding to remain after risk factor adjustment. Both of these are often not (or only partially) the case: In our UKB dataset, for instance, the risk factors BMI and uACR are available only from study center assessments, but not from eHRs; and in addition, it is very likely that further (unknown, unaccounted for) risk factors exist, which is why some degree of unmeasured confounding cannot be ruled out here. And while off-the-shelf bias correction methods for linear and binary outcomes exist [Dudbridge et al., 2019; Mahmoud et al., 2022; Donovan et al., 2025], their applicability to multi-state models is still to be explored.
Finally, as UKB participants are rather healthy [Fry et al., 2017], CKD progressions in this UKB dataset are relatively rare. This, in turn, makes the estimation of (especially small-scale, genetic) risk factors more difficult. It would thus be interesting to conduct similar CKD onset and progression analyses in less healthy cohorts.
@@
Multi-state PAMs (including all necessary preprocessing) are implemented in the R package pammtools [Bender and Scheipl, 2018], which uses mgcv::gam()/bam() [Wood and Wood, 2015] for efficient and flexible estimation. All code required for running the simulations (Section 4), for the UK Biobank analyses (Section 5), as well as for creating the corresponding figures and tables, is available from the GitHub repository https://github.com/survival-org/msm4diseaseHistories.
@@ A.1.1 Transformation to piecewise exponential data
Note that accounting for left-truncation, time-varying covariates, and time-varying effects all require similar data transformations and can therefore seamlessly be incorporated into this data preprocessing step [Piller et al., 2025]. To capture changes in time-varying covariates accurately, additional time cut points can be added at times where the time-varying covariates change [Bender and Scheipl, 2018].
PEMs then use $d_{ij}\sim Po(\mu_{ij})$ as outcome of a Poisson regression model with expected conditional event rate $\mu_{ij}=h_{ij}y_{ij}$ and logarithmized time-at-risk entering as offset $o_{ij}$, to estimate the hazard rate $h_{ij}$ as the product of a piecewise constant baseline hazard rate (for each interval) and the exponential of a linear predictor. PEMs can be shown to be equivalent to the Cox Proportional Hazards model under certain assumptions [Whitehead, 1980], but since they do not rely on the non-parametric estimation of the baseline hazard (due to assuming piecewise-constant hazards), they are fully parametric with respect to parameter estimation. Still, we here refer to PEMs as semi-parametric approach, analogously to the Cox model, because no parametric event time distribution is assumed - in contrast to parametric approaches such as Accelerated Failure Time models [Kalbfleisch and Prentice, 2002].
@@ A.2 Additional information on CKD onset and progression analyses in UK Biobank
A.2.1 Propensity score-based correction for index event bias
@@ A.4 Supplementary tables
Table A.3: Mean coverage of fixed effect log-hazards estimated using SSTS and MTS PAMs on data simulated from SSTS and MTS DGPs.